# Dexterous Manipulation through Imitation Learning: A Survey

### v3

@@ I Introduction
The main purpose of this survey is to provide an overview of IL-based dexterous manipulation approaches. The rest of this article is organized as follows: Section II presents an introduction to dexterous manipulation and IL in detail. Subsequently, we discuss the state-of-the-art IL-based dexterous manipulation techniques and highlight notable achievements in this field in Section III. Section IV discusses end-effectors for dexterous manipulation. Moreover, we discuss teleoperation systems and datasets in Section V VI. We summarize existing challenges and propose future directions for research in this rapidly evolving field in Sections VII and VIII, respectively. Finally, conclusions are made in Section IX. An overview of this survey is shown in Fig. 2. By synthesizing the existing body of knowledge, this survey aspires to serve as a valuable resource for researchers and practitioners seeking to advance the capabilities of robotic systems through the synergy of IL and dexterous manipulation.
@@ II-B Imitation Learning
Approach Key Characteristics Pros Cons
Behavioral Cloning [53, 54, 20, 55, 14, 56, 57, 29, 53, 26, 58, 59, 60, 32, 61, 62] Supervised learning paradigm. Direct mapping from states to actions. No reward signals or exploration. Simple and easy to implement. Data-efficient with large
@@
Computationally intensive. Non-unique reward solutions.
@@
Requiring hierarchy design. Requiring training coordination.
@@
Continual skill acquisition. Adapting previously learned behaviors. Flexible to evolving tasks. Reducing forgetting of old skills. Risk of catastrophic forgetting. Requiring ongoing expert input.
@@ III-C2 Research Progress
In addition to addressing the quality and availability of expert demonstrations, recent studies have sought to improve the performance of GAIL-based dexterous manipulation methods in other aspects. For instance, TRAIL [85] introduces constrained discriminator optimization to prevent the discriminator from focusing on spurious, task-irrelevant features such as visual distractors, thereby preserving meaningful reward signals and enhancing task performance. In the context of human imitation, Antotsiou et al. [86] combine inverse kinematics and particle swarm optimization with GAIL to mitigate sensor noise and domain discrepancies, enabling robots to autonomously grasp objects in simulation environments. Furthermore, P-GAIL [87] incorporates entropy-maximizing deep P-networks into GAIL to improve policy learning in deformable object manipulation tasks.
@@ III-D1 Description
The parameters of the high-level and low-level policies in HIL are typically determined through three approaches: (1) Learning from demonstrations, which utilizes expert demonstrations to train both levels of policies; (2) Optimization, which applies RL or other optimization methods to refine the policies; and (3) Manual tuning, which manually adjusts policies during the initial stages or for specific task requirements. A key advantage of HIL is its ability to reduce the complexity of direct action-space search by decomposing tasks into hierarchical structures. This decomposition not only improves learning efficiency, particularly in long-horizon tasks but also enhances generalization and task success rates by enabling optimization at multiple levels.
@@ III-D2 Research Progress
In recent years, HIL has achieved significant progress in robotics, particularly in task decomposition and skill generalization. CompILE [88] enhances generalization in complex environments by decomposing tasks into independent sub-tasks. This compositional approach sets the foundation for subsequent research, particularly for tasks involving long temporal sequences. Building on this, Xie et al.[89] apply HIL to dual-arm manipulation and introduce the HDR-IL framework, which decomposes tasks into multiple motion primitives and employs graph neural networks to model object relationships.
HIL has also been successfully applied to various dexterous manipulation tasks. ARCH, proposed by Sun et al.[90], presents a framework tailored for long-horizon, contact-rich robotic assembly. It combines a low-level library of predefined skills with IL for high-level policy, enabling efficient composition and adaptation of skills to handle complex, high-precision tasks. XSkill, introduced by Xu et al.[91], further extends HIL to cross-modal skill discovery. Their framework demonstrates how learning skill mappings across different modalities allow robots to generalize to diverse environments. Additionally, Wan et al. [92] propose LOTUS, which focuses on maintaining skill continuity and stability in dynamic environments by decomposing tasks into continuous sub-tasks and enabling policy adaptation across varying conditions. To efficiently train both high-level and low-level policies, recent works have explored the use of play data. Wang et al.[93] propose MimicPlay, which leverages unstructured human play interactions to guide robot manipulation. This approach learns high-level latent plans from play data and uses them to train a low-level visuomotor controller with only a small amount of demonstrations. Similarly, Lin et al.[94] introduce H2RIL, which utilizes cross-domain demonstrations by extracting interaction-aware skill embeddings from task-agnostic play data. These embeddings are aligned with human videos via temporal sequence contrastive learning, enabling the system to generalize to novel, composable tasks and adapt to out-of-distribution scenarios.
@@ III-D3 Discussion
In summary, HIL has demonstrated significant advantages in task decomposition, skill generalization, and handling long-horizon tasks. By introducing hierarchical structures and multi-level control strategies, substantial progress has been made across various dexterous manipulation tasks. However, current research struggles with achieving adaptability in cross-modal skill generalization and ensuring model robustness and continuity in dynamic environments. In particular, when task environments change substantially, rapidly adapting and optimizing existing skill libraries to enable smooth task transitions remains an open problem.
@@ III-E1 Description
Continual Imitation Learning (CIL) is an approach that integrates continual learning with IL, aiming to enable agents to continuously acquire and adapt skills by imitating expert behaviors in dynamically changing environments. Specifically, in the initial phase, the agent learns fundamental skills from expert demonstrations. In subsequent phases, the agent incrementally accumulates knowledge, adapts to new tasks or environments, and mitigates the risk of forgetting previously acquired skills.
@@ III-E2 Research Progress
CIL has emerged as a key research direction in robotics, with the goal of enabling robots to continually acquire new skills from task demonstrations while mitigating catastrophic forgetting of previously learned tasks. To this end, various approaches have been especially proposed for dexterous manipulation.
Early studies in CIL focus on enabling robots to switch between tasks without compromising previously acquired skills, laying the groundwork for subsequent research [95]. However, these approaches often require substantial storage and computational resources, limiting their practicality in real-world applications. To address these limitations, researchers propose task-specific adapter structures, introducing lightweight adapters to facilitate seamless task switching [96]. This method effectively reduces storage overhead and improves adaptability to new tasks, though its performance tends to decline when tasks differ significantly. Further advancements explore the use of unsupervised skill discovery to enhance adaptability [92]. By dynamically generating new skills and integrating them into the robot’s existing skill set, this approach improves the robot’s ability to handle complex and evolving tasks. While promising results have been demonstrated in simulation, the effectiveness and generalization of these skills in real-world scenarios remain to be validated.
Further advancements in CIL have introduced the concept of learning unified policies through behavior distillation [97]. Unlike earlier works, this approach addresses the challenge of multi-task learning by employing a single policy structure, thereby eliminating the need for additional adapters when new tasks are introduced. However, designing such unified policies remains challenging, particularly in maintaining performance on previously learned tasks while accommodating new ones. In parallel, the emergence of GAN has inspired the development of CIL-based Deep Generative Replay (DGR)[98], which enables robots to continually acquire new skills by generating synthetic task trajectories. This approach alleviates the need to store past task data or recreate previous environments, effectively reducing storage overhead. Nevertheless, ensuring the quality and consistency of the generated trajectories remains an open challenge. Additionally, self-supervised learning has been explored to extract skill abstractions and expand the applicability of CIL to complex control tasks[99]. This method demonstrates that, even in the absence of explicit task demonstrations, robots can continually acquire new skills through self-supervision.
@@ III-E3 Discussion
In summary, current research in CIL for dexterous manipulation mainly focuses on effective multi-task learning, the application of DGR techniques, and self-supervised skill abstraction. Through these methods, researchers have addressed some of the core challenges of continual learning, particularly in task switching and skill integration. However, significant challenges remain for practical deployment. The quality and consistency of generated data are not yet optimal, which may impact the accuracy of decision-making and task execution. Additionally, although some approaches have reduced storage and computational demands, resource consumption remains a limiting factor in more complex scenarios. Furthermore, the generalization capabilities of current methods, particularly in handling diverse tasks and dynamic environments, are still insufficient for real-world applications.
IV End Effectors for Dexterous Manipulation

### v4

@@ Abstract
Note to Practitioners
This work explores the intersection of IL and dexterous manipulation. With promising applications in manufacturing, healthcare, and home robotics, IL-based approaches allow robots to handle delicate objects, execute precise tasks, and operate in diverse, unstructured environments. However, key challenges remain—particularly in collecting high-quality demonstrations and enabling generalization from limited data, both of which are critical for realworld deployment. This survey aims to serve as a practical and accessible guide for practitioners seeking to apply IL-based methods to develop more capable and adaptable robotic systems in real-world scenarios.
@@ I Introduction
Building on significant advancements in the intersection of imitation learning (IL) and dexterous manipulation, several recent surveys have explored key aspects of this rapidly evolving field. Zare et al. [36] provide an extensive overview of the applications, progress, and challenges of imitation learning in both robotics and artificial intelligence. Arora and Doshi [37] focus on inverse reinforcement learning (IRL), particularly the process of inferring reward functions from observed behavior. Han et al. [38] delve into deep reinforcement learning (DRL), reviewing recent methods designed to optimize DRL algorithms for real-world applications. Li et al. [39] offer a comprehensive review of data collection methods and skill learning frameworks for robotic dexterous manipulation, emphasizing the key challenges in this area. Pitkevich and Makarov [40] address the Sim-to-Real transfer problem in DRL, discussing the deployment of robots trained in simulation to real-world environments. Welte and Rayyes [41] examine the potential of interactive imitation learning for humanoid robots, with a focus on how these methods can be transplanted for dexterous tasks. Finally, Tsuji et al. [42] provide a systematic review of imitation learning techniques for contact-rich tasks, discussing the unique challenges and emerging trends in this field. In contrast to these existing works, which primarily focus on specific aspects of imitation learning or dexterous manipulation, this survey aims to provide a comprehensive overview of IL-based dexterous manipulation approaches.
The rest of this article is organized as follows: Section II presents an introduction to dexterous manipulation and IL in detail. Subsequently, we discuss the state-of-the-art IL-based dexterous manipulation techniques and highlight notable achievements in this field in Section III. Section IV discusses end-effectors for dexterous manipulation. Moreover, we discuss teleoperation systems and datasets in Section V. We summarize existing challenges and propose future directions for research in this rapidly evolving field in Section VI. Finally, conclusions are made in Section VII. An overview of this survey is shown in Fig. 2. By synthesizing the existing body of knowledge, this survey aspires to serve as a valuable resource for researchers and practitioners seeking to advance the capabilities of robotic systems through the synergy of IL and dexterous manipulation.
@@ II-B Imitation Learning
II-C Connections to Human Motor Learning
IL is fundamentally motivated by the way humans acquire motor skills through observation, demonstration, and practice. From a cognitive perspective, Bandura’s social learning theory [60] emphasizes that behavior can be learned by observing others, rather than solely through reinforcement or trial-and-error. This insight has inspired IL paradigms that rely on demonstration-based policy generation instead of environment-driven exploration.
At the neural level, the discovery of mirror neurons in primates provides biological evidence that action observation and execution may share overlapping representations [61]. These findings suggest that learning by watching others perform a task can activate internal motor plans, a mechanism analogous to behavioral cloning in robotics. Furthermore, models of human motor control, such as the internal model theory [62] and optimal feedback control [63], offer insights into how the brain predicts, plans, and corrects motor behaviors. These principles have influenced the design of IL algorithms that incorporate predictive models or closed-loop feedback refinement. For example, Nah et al. [64] present a modular robot control framework using Dynamic Movement Primitives (DMP) and Elementary Dynamic Actions (EDA), employing dynamical systems to predict and correct motor behaviors in a manner similar to internal models in the brain.
In summary, the foundations of IL are deeply rooted in cognitive and neuroscientific principles, from observational learning and mirror neuron systems to predictive motor control and hierarchical skill organization. These biological insights continue to shape IL algorithms, guiding the development of methods that mimic the way humans learn and adapt motor behaviors.
II-D Long-horizon Dexterous Manipulation
Long-horizon dexterous manipulation involves executing complex tasks that require both fine-grained control and long-term planning. These tasks typically span long time horizons and require coherent, efficient execution of sequential actions. Without appropriate structure, such tasks can overwhelm agents, complicating both planning and execution. Task decomposition addresses this challenge by enabling agents to focus on different task components at various levels of abstraction, thereby facilitating the management of long-horizon manipulation [65]. In the context of IL, task decomposition is commonly realized through temporal abstraction, often implemented via techniques such as option discovery [66] or hierarchical policy structures [67, 68].
Temporal abstraction enables an agent to make decisions through temporally extended action sequences, thereby accelerating learning and simplifying complex tasks. Option discovery refers to the automated process of identifying and constructing these temporal abstractions, known as options, that comprise an internal low-level policy, an initiation set, and a termination condition. The options framework [69, 70] generalizes the traditional action concept by treating options as temporally extended actions, allowing planning over multiple time scales. The option-critic architecture [71] and its variants are widely adopted to automatically extract options from demonstration data, which are then scheduled by high-level policies to effectively organize the temporal structure of tasks. For example, DDCO [72] autonomously discovers and learns hierarchical options within continuous action spaces from expert demonstrations, enabling unsupervised task decomposition and hierarchical imitation learning for robotic control.
Hierarchical reinforcement learning is often integrated with IL to tackle long-horizon tasks and complex policy learning [73]. The fundamental principle involves leveraging human demonstrations to decompose tasks along temporal or functional dimensions, which reduces the policy search space and enhances both learning efficiency and generalization. For example, SRT-H [74] proposes a hierarchical language-conditioned IL framework that integrates high-level instruction parsing with low-level dexterous control to allow the execution of end-to-end autonomous surgical tasks.
II-E Multi-agent or Collaborative Dexterous Manipulation
Multi-agent or collaborative dexterous manipulation refers to the system-level problem in which two or more agents endowed with multi-fingered dexterous manipulation capabilities, such as robot arms, dexterous hands, or mobile manipulators, jointly accomplish a high-dimensional, long-horizon, contact-rich task. The scenario emphasizes:
  • heterogeneous or homogeneous agents sharing workspace and task objectives;
  • explicit or implicit communication (force, vision, language instructions) for action synchronization and load distribution;
  • leveraging the redundant degrees of freedom of multiple arms/hands to realize complex operations unattainable by a single arm (e.g., bimanual valve turning, multi-human cooperative suturing).
To enable collaborative dexterous manipulation, Kim et al. [74] introduced the SRT-H framework, which employs a hierarchical policy integrating high-level language instructions with low-level trajectory imitation. In a dual-arm da Vinci suturing task, the framework achieves precise bimanual coordination over long horizons, relying solely on human demonstrations without any reward supervision. Building on this line of research, BUDS [21] introduced a “stabilize-then-act” role decomposition with a visual keypoint stabilization mechanism, attaining 76.9% task success on complex bimanual dexterous tasks without reward signals or RL fine-tuning. Extending further, Bi-DexHands [75] framed bimanual dexterous manipulation as a comprehensive suite of over 20 cooperative sub-tasks, providing the first large-scale benchmark for evaluating the scalability of multi-agent RL in physically coordinated manipulation. More recently, BiDexHD [76] advanced collaborative dexterous manipulation by integrating multi-task human demonstration data with teacher–student policy learning. This framework enables robots to efficiently acquire a wide range of cooperative bimanual skills across numerous auto-constructed tasks, substantially improving task completion rates and zero-shot generalization.
In general, the core challenge of multi-agent or collaborative dexterous manipulation lies in learning generalizable cooperative policies within large-scale continuous control spaces while satisfying safety, real-time responsiveness, and consistency with human demonstrations, particularly under the imitation-learning paradigm.
@@ II-E Multi-agent or Collaborative Dexterous Manipulation
Approach Key Characteristics Pros Cons Key Applications
Behavioral Cloning [77, 78, 20, 79, 14, 80, 81, 29, 77, 26, 82, 83, 84, 32, 85, 86] Supervised learning paradigm. Mapping from states to actions. No reward signals or exploration. Simple and easy to implement. Data-efficient with large
@@
Grasping and pick-and-place tasks. Short-horizon, structured tasks with sufficient data.
@@
Computationally intensive. Non-unique reward solutions. Complex tool-use tasks. Assembly tasks with implicit objectives.
@@
Long-horizon dexterous tasks. In-hand manipulation with sparse demonstrations.
@@
Requiring hierarchy design. Requiring training coordination. Multi-step assembly and contact-rich tasks. Skill chaining and long-horizon manipulation.
@@
Continual skill acquisition. Adapting previously learned behaviors. Flexible to evolving tasks. Reducing forgetting of old skills. Risk of catastrophic forgetting. Requiring ongoing expert input. Lifelong learning for multi-task manipulation. Adapting to changing tools or objects.
@@ III-A2 Research Progress
Compared to variational autoencoders (VAEs) [123] and generative adversarial networks (GANs) [124], diffusion models offer several advantages for imitation learning, particularly in modeling complex, multimodal action distributions [32, 125]. Chen et al. [84] proposed using diffusion models to augment behavioral cloning and conducted a comprehensive comparison with other generative models, including VAEs, GANs, and EBMs(energy-based models)/implicit models, across different implementations in terms of model architecture, sampling/inference cost, and performance on various continuous control and manipulation tasks. Experimental results demonstrated that diffusion-augmented behavioral cloning outperformed standard BC and several other generative model variants on multiple tasks.
@@ III-C2 Research Progress
In addition to addressing the quality and availability of expert demonstrations, recent studies have sought to improve the performance of GAIL-based dexterous manipulation methods in other aspects. For instance, TRAIL [107] introduces constrained discriminator optimization to prevent the discriminator from focusing on spurious, task-irrelevant features such as visual distractors, thereby preserving meaningful reward signals and enhancing task performance. In the context of human imitation, Antotsiou et al. [Antotsiou2018TaskOrientedHM] combine inverse kinematics and particle swarm optimization with GAIL to mitigate sensor noise and domain discrepancies, enabling robots to autonomously grasp objects in simulation environments. Furthermore, P-GAIL [108] incorporates entropy-maximizing deep P-networks into GAIL to improve policy learning in deformable object manipulation tasks.
@@ III-D1 Description
The parameters of the high-level and low-level policies in HIL are typically determined through three approaches: (1) Learning from demonstrations, which utilizes expert demonstrations to train both levels of policies; (2) Optimization, which applies RL or other optimization methods to refine the policies; and (3) Manual tuning, which manually adjusts policies during the initial stages or for specific task requirements. A key advantage of HIL is its ability to reduce the complexity of direct action-space search by decomposing tasks into hierarchical structures. This decomposition not only improves learning efficiency, particularly in long-horizon tasks, but also enhances generalization and task success rates by enabling optimization at multiple levels.
@@ III-D2 Research Progress
In recent years, HIL has made significant progress in task decomposition and skill generalization. CompILE [109] enhances generalization in complex environments by decomposing tasks into independent sub-tasks, laying the foundation for subsequent work, particularly for tasks with long temporal dependencies. In dexterous manipulation, ARCH [111] integrates a low-level library of predefined skills with a high-level IL policy, enabling efficient skill composition and adaptation for complex, high-precision tasks. Meanwhile, Wan et al. [113] emphasize maintaining skill continuity and stability in dynamic environments by decomposing tasks into continuous sub-tasks and enabling adaptive policy transitions under varying conditions.
To enhance scalability, recent research has increasingly explored automatic hierarchy discovery to reduce reliance on manually designed skill structures. Unsupervised skill segmentation methods have shown strong potential for identifying reusable sub-skills and task boundaries directly from raw data. For example, Gehring et al.[127] propose hierarchical exploration strategies that use unsupervised clustering to learn multi-level reusable sub-policies. More recently, Jansonnie et al.[128] present a skill discovery framework that leverages automatic task generation and asymmetric self-play to produce diverse manipulatory behaviors, which are then embedded into reusable primitives suitable for hierarchical policies. Similarly, contrastive learning approaches [129] discover skills by maximizing mutual information across trajectory embeddings, resulting in more robust and generalizable skill representations.
From a data-centric perspective, play data has been utilized to train both high-level and low-level policies. Wang et al.[114] propose MimicPlay, which exploits unstructured human play interactions to learn high-level latent plans and trains a low-level visuomotor controller using only a small number of demonstrations. Likewise, Lin et al.[115] introduce H2RIL, which extracts interaction-aware skill embeddings from task-agnostic play data and aligns them with human videos via temporal contrastive learning, enabling generalization to novel, composable tasks and adaptation to out-of-distribution scenarios.s
@@ III-D3 Discussion
In summary, HIL has demonstrated significant advantages in task decomposition, skill generalization, and handling long-horizon tasks. By introducing hierarchical structures and multi-level control strategies, substantial progress has been made across various dexterous manipulation tasks. However, most current approaches still depend on manually designed hierarchies, which limits autonomy in unstructured or dynamic environments. Automatic hierarchy discovery methods, such as unsupervised skill segmentation or contrastive learning-based sub-task discovery, are essential for enabling robots to autonomously identify reusable skills and adapt their hierarchy to new tasks. Furthermore, ensuring robustness and continuity when skill libraries need to be updated or extended, particularly under changing task conditions, remains an open problem.
@@ III-E1 Description
Continual Imitation Learning (CIL) combines continual learning with IL to enable robots to continuously acquire and adapt new skills from expert demonstrations while avoiding the forgetting of previously learned knowledge when adapting to new tasks in dynamic environments. Specifically, in the initial phase, the agent learns fundamental skills from expert demonstrations. In subsequent phases, the agent incrementally accumulates knowledge, adapts to new tasks or environments, and mitigates the risk of forgetting previously acquired skills.
@@ III-E2 Research Progress
Various CIL approaches have been developed for dexterous manipulation, each aiming to balance memory retention and adaptation. Early studies [116] relied on storing large amounts of task data, leading to high storage and computational costs. To address this, task-specific adapter structures introduced lightweight, modular components for seamless task switching [117], though their performance declines when task variations are substantial.
To address catastrophic forgetting more systematically, regularization-based methods[130] have gained prominence. For instance, Elastic Weight Consolidation (EWC)[131] penalizes large updates to parameters critical for previously learned tasks by incorporating Fisher information into the loss function. Similarly, knowledge distillation-based approaches[132] preserve prior skills by encouraging new policies to mimic outputs from older policies. Rehearsal-based strategies offer another solution, such as experience replay, where past experiences or generated trajectories are interleaved with new task data to reduce forgetting[133]. Generative models, such as Deep Generative Replay (DGR), further alleviate storage constraints by synthesizing realistic task trajectories instead of storing raw data [119], although ensuring trajectory fidelity remains challenging.
Beyond these, unsupervised skill discovery has been explored to improve adaptability and generalization [113]. It dynamically generates new skills and integrates them into the robot’s repertoire, but has yet to demonstrate robust real-world generalization. Unified policy learning via behavior distillation [118] provides an alternative by employing a single shared policy across tasks, eliminating the need for task-specific modules, though at the cost of increased task interference. Lastly, self-supervised learning techniques have shown promise in acquiring transferable skill abstractions without explicit demonstrations  [120].
@@ III-E3 Discussion
In summary, CIL for dexterous manipulation has explored a range of techniques, such as adapter-based methods, regularization approaches, knowledge distillation, rehearsal strategies, and generative models. These approaches collectively aim to mitigate catastrophic forgetting and enhance task-switching efficiency, but significant challenges remain. The effectiveness of regularization and knowledge distillation often depends on carefully balancing new and old task performance, which becomes difficult when tasks differ substantially. Similarly, the stability and performance of rehearsal or generative replay methods are highly sensitive to the quality and diversity of stored or synthesized data. Although modular adapters and regularization techniques have reduced memory demands, computational overhead and scalability continue to pose barriers for real-world deployment.
TABLE II: A Comparison of IL Methods in Cost, Feasibility, Efficiency, and Complexity
Category Training Cost Inference Latency Sample Efficiency Computational Complexity Real-Time Feasibility Notes
BC Low Very Low (<1<1 ms) Low–Moderate (requires many demos; improves with online correction) Low Highly suitable for high-DOF real-time control; main bottleneck is data quality and covariate shift.
IRL High Low–Moderate (1–5 ms) Moderate–Low High Training is computationally expensive; inference feasible but may require reward computation if deployed online.
GAIL High Low–Moderate (2–5 ms) Moderate High Inference lightweight once trained; training unsuitable for strict real-time constraints.
HIL Moderate–High Moderate (2–10 ms) High (can reuse sub-policies) Moderate–High Two-tier policy execution increases latency; optimization needed for tight loops.
CIL High Moderate (2–8 ms) Moderate–High Moderate–High Online adaptation may break hard real-time constraints; feasible if adaptation is offloaded.
TABLE III: A Comparison of IL Methods by Optimization and Convergence
Method Key Hyperparameters Common Optimizers Regularization Strategies Convergence Behavior Notes
BC Learning rate, batch size, network depth Adam, RMSprop L2 weight decay, dropout Converges quickly; risk of overfitting on small datasets; performance limited by covariate shift.
IRL Reward learning rate, policy learning rate, discriminator updates Adam for policy & reward networks Entropy regularization, reward clipping Slower convergence due to reward inference; reward shaping essential for stability.
GAIL Policy LR, discriminator LR, batch size, rollout length Adam (policy & discriminator) Gradient penalty, spectral norm Sensitive to discriminator–policy update ratio; prone to oscillation without tuning.
HIL Subpolicy LR, meta-controller LR, option horizon Adam (separate for high- & low-level) Option dropout, inter-option regularization Convergence depends on option discovery quality; mitigates long-horizon credit assignment.
CIL Base LR, regularization strength for knowledge retention, replay buffer size Adam, SGD with momentum Elastic Weight Consolidation (EWC), replay-based regularization Trains indefinitely with periodic adaptation; stability–plasticity trade-off crucial.
III-F Comparative Analysis of Imitation Learning Approaches
While Table I outlines the general pros and cons of different IL approaches, the challenges inherent to dexterous manipulation amplify these limitations and deserve deeper analysis. Table II summarizes the trade-offs between efficacy and computational complexity across imitation learning paradigms. BC offers a simple, data-efficient solution but is highly vulnerable to distribution shifts, where small deviations can rapidly compound into task failure. IRL addresses this by inferring expert reward functions to improve generalization, though designing meaningful rewards for complex tasks remains difficult and computationally intensive. GAIL combines the strengths of BC and IRL by adversarially matching expert behavior without explicit reward design, yet it often suffers from instability and mode collapse in multi-modal tasks. Structurally, Hierarchical IL builds on these methods by decomposing tasks into modular skill primitives, improving scalability and reusability; however, reliance on hand-crafted hierarchies limits adaptability to novel scenarios. Continual IL addresses the evolving nature of manipulation tasks by supporting incremental skill acquisition, but dynamic contact interactions increase the risk of catastrophic forgetting. In summary, BC is well-suited for short-horizon, data-rich settings, while IRL and GAIL provide more robust generalization when expert intent must be inferred. Hierarchical and continual strategies extend these foundations to support long-term skill composition and adaptation. As detailed in Table III, the optimization specifics and convergence behavior of imitation learning methods are analyzed.
Hybrid strategies hold promise for achieving both robustness and scalability in complex, long-horizon tasks by integrating complementary strengths of different paradigms. A key direction involves integrating hierarchical planning with adversarial imitation by employing a high-level policy that generates symbolic or language-conditioned subgoals, while the low-level policy is optimized through adversarial learning to match expert trajectories under complex contact dynamics. This coupling allows the high-level controller to focus on long-horizon task composition while the low-level controller benefits from the sample efficiency and robustness of adversarial training in modeling fine-grained manipulation primitives. Another approach is embedding continual learning mechanisms into hierarchical architectures, enabling incremental skill acquisition while preserving previously learned behaviors through replay-based regularization or generative memory. Reward shaping derived from IRL can be combined with supervised BC objectives to improve generalization and reduce compounding errors in data-limited settings. Furthermore, hybrid systems can integrate meta-learning for rapid skill adaptation, while lightweight model-based modules, such as MPC for local corrections, complement model-free IL for enhanced stability and precision.
An increasingly important aspect of hybrid IL strategies is the integration of model-based control with imitation learning, which is often underexplored. Model-based components, such as dynamics models or physics priors, can provide structured guidance for low-level policy optimization, reducing reliance on dense supervision or large datasets. For instance, MPC can be used to plan feasible trajectories or correct drift during execution, while the IL policy handles task-level reasoning and adaptation. Conversely, IL can help bootstrap model-based controllers in regions where dynamic models are inaccurate or data is sparse. This synergistic integration allows systems to exploit the strengths of both paradigms: the interpretability, adaptability, and safety of model-based control with the expressiveness and flexibility of data-driven IL. In real-world manipulation, where uncertainties and contact-rich dynamics are prevalent, such hybrid control strategies have shown promise in improving generalization, robustness, and sample efficiency.
TABLE IV: Comparative Performance of Three End-Effector Types in Dexterous Manipulation
End-Effector Type DoF Dexterity & Design Notes Actuation Principle Control Precision Adaptability Cost Typical Applications Suitability for IL
Two-Fingered Gripper 1–2 Simple parallel jaws; limited in-hand re-configuration Single-motor, linkage or direct-drive Medium Low Low Pick-and-place, repetitive tasks Low data demand, easy to train, strong generalization
Three-Fingered Robotic Claw 6–9 Three opposed fingers (Fig.4); hybrid under-actuation yields cylindrical/spherical grasps with modest dexterity Tendon-driven or linkage + spring coupling Medium–High Medium Medium General grasping, limited in-hand manipulation Moderate data demand, balances flexibility and training cost
Multi-Fingered Anthropomorphic Hand 15–25 Full five-finger layout (Fig. 3); high DoF enables human-like precision and fingertip control Remote tendon, direct-drive, or hybrid transmissions High High High Medical procedures, precision assembly, complex tasks High data demand, expensive training, limited generalization
IV End-effectors for Dexterous Manipulation